OpenStudy (jbeanz97):

can someone help with an algebra1 question? In 2005, a Nevada silver mine exported 200 tons of silver. In 2008, as the silver supply became exhausted, it only exported 25 tons of silver. What is the rate of change of the mine’s silver output?

OpenStudy (jbeanz97):

\[A. -\frac{ 175 }{ 3 } B.-\frac{ 1 }{ 50 } C.\frac{ 175 }{ 3 } D.75\]

OpenStudy (jbeanz97):

@ganeshie8 @phi @mathstudent55 help?